# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v3

@@ 4.2 Model Comparison
Table,2, 3, 4, 5 demonstrate the comparison results on ADNI, TaoWu, PPMI, and ABIDE datasets. The mean and standard deviation of ten-fold cross-validation were listed, with the best results highlighted in bold and the second-best results in underlining. Compared with traditional ML methods, general-purposed GNN-based models for brain network do not show a significant advantage and attain similar performance in most cases. This finding is consistent with the results reported in existing papers. This is because, under the same input conditions, neural network models typically have a larger number of parameters than conventional machine learning methods when utilizing edge features for better interpretability, making them more susceptible to overfitting. Transformer-based models predominantly outperform GNN-based models across all four datasets in most scenarios, highlighting their superior capability to capture the global structure of brain connectomes. HGNN-based models generally do not demonstrate superiority over other models, with notable performance only observed on the PPMI dataset. This phenomenon can be attributed to two main reasons: first, such models introduce excessive parameters, posing a risk of overfitting; second, their k-hop neighbor-based modeling approach fails to effectively capture the co-fluctuation characteristics among multiple brain regions, thereby inadequately representing HOIs. In PH-based models, such as ATPGCN and Brain-HORs, which introduce higher-order loops or connected component features extracted via persistent homology for brain disease diagnosis, demonstrate significant superiority compared to both graph neural network (GNN)-based models and traditional machine learning models that rely solely on edge features. This highlights the necessity of incorporating higher-order topological features in brain disease diagnosis methods.
@@
Our HOI-brain outperforms all 20 baselines on all datasets. In particular, it can be found that the proposed HOI-brain outperforms the traditional machine learning models on all datasets, with average accuracy improved by 12.5 (%), average precision improved by 12.9(%), average recall improved by 14.3(%) and F1-score improved by 12.7(%), respectively. And the proposed HOI-brain outperforms the GNN-based models on all datasets, with average accuracy improved by 19.6(%), average precision improved by 22.9(%), average recall improved by 19.5(%), and F1-score improved by 19.5(%), respectively. This indicates that our model, by incorporating higher-order topological features, is more beneficial for brain disease diagnosis compared to using only low-order edge features. Furthermore, the proposed HOI-Brain outperforms Transformer-based models on all datasets, with average accuracy improved by 13.0(%), average precision improved by 13.4(%), average recall improved by 10.3(%), and F1-score improved by 10.9(%), respectively. This highlights that our model, through the multi-channel brain network Transformer, can effectively and holistically integrate complementary information from heterogeneous higher-order and low-order topological features while preserving channel-specific properties. Next, when compared to HGNN-based models, HOI-Brain achieves remarkable improvements with average accuracy improved by 21.8(%), precision improved by 27.0(%), recall improved by 19.4(%), and F1-score improved by 20.5(%). These results clearly demonstrate that our proposed HOI-Brain framework significantly advances the modeling of HOIs through the Multiplication of Temporal Derivatives and the extraction of topological features using persistent homology in brain networks. When specifically compared to Persistent Homology (PH) based models, HOI-Brain achieves significant improvements with average accuracy increased by 17.4(%), precision enhanced by 21.4(%), recall boosted by 14.5(%), and F1-score elevated by 16.2(%). This is partly attributed to our method linking signed HOIs in the brain to higher-order organizations, and partly due to our exploration of more complex signed quadruplet structures and voids (2D holes) formed by triangular faces and quadruplet.
@@ 4.3.1 Ablation study
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of different components on overall performance. Our first aim was to evaluate the effectiveness of more complex signed quadriplet structures and voids (2D holes) formed by triangular faces. Specifically, following the method in (Santoro et al., 2024), we extract lower-order edge features, violating triangles, one-dimensional loops. In addition, we extract unsigned good quadruplets and unsigned two-dimensional voids based our methods. We implement it within different combinations of these features on all datasets. The experimental results presented in Table 6, 7, 8, 9, indicate that integrating both higher-order and lower-order interactions proves to be more effective for brain disorder diagnosis compared to using lower-order interactions alone. Furthermore, quadruplet interactions captured by HOI-Brain have been demonstrated to be more effective in improving diagnosis than triplet interactions. This may be due to the fact that a significant proportion of triplet interactions can be decomposed into a linear combination of pairwise interactions, which is insufficient to capture the true HOIs. Additionally, distinguishing between positively and negatively synergistic HOIs to generate signed higher-order structures, as opposed to unsigned ones, can further improve the effectiveness of brain disease diagnosis, which may due to the fact that it can offer detailed insights into the communication within the brain.
Our second aim was to evaluate the effectiveness of different components in multi-channel brain network Transformer. The different variants are: not adding signed higher-order features decoupling mechanism (wo-signed), not adding attention-guided feature fusion mechanism (wo-fusion), and not adding orthonormal clustering readout of multiple channels(wo-cluster). From Fig. 4, we can see that adding signed higher-order features decoupling mechanism and attention-guided feature fusion mechanism are always helpful as HOI-brain wins all the cases over the variant (wo-signed) or (wo-fusion). This result is consistent with our assumption that, by introducing the attention mechanism, our model can adaptively fuse positive and negative information as well as information from different channels, thereby emphasizing discriminative features while suppressing redundant information. Incorporating orthonormal clustering readout across multiple channels (wo-cluster) is highly beneficial, as this approach effectively captures both low-level and high-level feature patterns, which often reside within distinct functional modular structures.
@@ 4.3.2 Hyperparameter analysis
To further demonstrate how the design of orthonormal clustering readout across multiple channels affects our model’s performance, we investigate a key hyperparameter: the number of clusters K. Specifically, we evaluate the method with K set to 2, 3, 4, 5, 10, 20, 50, and 100. The results of accuracy metric tuning this hyperparameter across all datasets are presented in Fig. 5. The results of other metrics are presented in Fig. 18. We observe that the model’s performance improves as K increases from 2 to 10 or 20 but declines when K rises from 10 or 20 to 100. This suggests that the optimal number of clusters is relatively small, reducing computational cost while aligning with the fact that the typical number of functional modules is fewer than 25, a finding consistent with previous studies (Kan et al., 2022b).
@@ 4.4.1 In-depth Analysis of Attention Mechanism
To better understand the mechanism of our model, which focuses on the most discriminative information in the brain for the diagnosis of brain diseases, we first performed a visualization analysis of the attention scores , of the attention-guided feature fusion mechanism for all patients of each type of disease, using ADNI and ABIDE as examples. Fig. 6 shows that our model is capable of adaptively learning weights according to the characteristics of different diseases, thereby integrating higher-order and lower-order topological information for the diagnosis of brain diseases. In addition, we observed that quadruplet-level interaction signatures are significantly more important than two-dimensional void descriptors and edge features, with their average attention scores exceeding 0.5 on both datasets.
@@
Next, we also conducted a visualization analysis of the attention scores , , , derived from the signed higher-order feature decoupling mechanism for all patients across each disease type, using ADNI and ABIDE as examples. Fig. 7 shows that across both datasets, the importance of negatively synergistic quadruplets is generally greater than that of positively synergistic information in most cases, while the importance of positively synergistic voids exceeds that of negatively synergistic information. This new phenomenon may provide a clearer direction for research into the pathological mechanisms of brain diseases.
@@
In order to conduct a more in-depth investigation into the aforementioned phenomena in the attention mechanism, inter-group comparisons were conducted of the signed higher-order topological features, the quadruplet-level interaction signatures and two-dimensional void descriptors. These were captured by the model in both the patient and healthy control groups. The ADNI dataset was used as an example. As demonstrated in Fig. 8, which presents a comparison of group differences in higher-order topological features, it was observed that the group differences in the quadruplet-level interaction signatures were found to be highly significant, with all statistical significance levels at compared to void descriptors. This finding is consistent with the results of our model visualisation. This phenomenon may be attributed to the disruption of HOIs within the brain during the progression of the disease, which in turn exerts an influence on the high-dimensional hole structures present within the brain. Furthermore, as the disease progresses, the number of positive quadruplets exhibits a gradual decrease, while the number of negative quadruplets initially decreases and subsequently increases. This finding suggests that during the transition from CN to MCI, there is a decline in synergistic interactions in the brain, whether positive or negative. Conversely, as the condition progresses from MCI to AD, brain regions have been observed to exhibit negatively synergistic interactions, concurrently reducing their excitability to preserve the functionality of diverse neural modules. With regard to the number of voids, although no clear trend was observed in the progression of the disease, it was found that positive voids exhibited more significant inter-group differences compared to negative voids. This finding lends further support to the hypothesis that they exhibit higher levels of attention scores. Consequently, CN appears to utilise positive information to construct the whole-brain topological structure, whereas AD relies more heavily on negative information. These findings may offer a novel perspective on the pathological mechanisms underlying Alzheimer’s disease.
@@ 4.4.2 In-depth Analysis of Cluster Mechanism
To better understand how our model leverages modular-level similarities between ROIs across different brain network patterns, we first provide a visualization of three attention matrices at the first layer of the multi-head self-attention module on the ADNI and ABIDE datasets in Fig. 9, 10. The attention scores are averaged across all subjects in the ADNI or ABIDE test set. This figure demonstrates that the learned attention scores reveal varying degrees of modularity in both low-order and higher-order patterns, highlighting the interpretability of our model. Furthermore, we provide a visualization of three cluster assignment matrices on the ADNI or ABIDE dataset in Fig. 11. The cluster assignment matrices are multiplied across layers and averaged across all subjects in the ADNI and ABIDE test set. Each row corresponds to an ROI and each column is a cluster. We observe that the assignment matrix, which is calculated based on the similarity between nodes and clusters, contains only one high-value entry per row in most cases, despite the fact that we did not impose any regularization constraints on the assignment matrix. This indicates that our model can effectively partition nodes into functional modules. Additionally, there are differences between the assignment matrices of low-order and higher-order patterns, which further demonstrates that varying degrees of modularity in both low-order and higher-order patterns. This result aligns with the visualization of three attention matrices, further indicating the interpretability of our model.
@@ 4.4.3 The important brain regions and interactions of brain regions
In this section, we further investigate whether the learned ROIs attention scores are interpretable and consistent with previous findings. Here, we take the ADNI, PPMI and ABIDE datasets as examples for discussion of three diseases. These three datasets contain larger sample sizes, enabling the reflection of more generalizable phenomena for brain diseases. First, we analyze the important brain regions and interactions of brain regions associated with brain diseases using HOI-Brain from a more comprehensive perspective rather than a low-level perspective as in previous studies (Kan et al., 2022b; Yu et al., 2024). For each subject, HOI-brain gets three attention matrixes at the first layer of multi-head self attention module. Higher attention scores indicate greater discriminative power of the regions for AD, PD, and ASD. For each disease, we obtain a comprehensive attention matrix that simultaneously captures both low-level and high-level patterns by performing a weighted summation of the three attention matrices for each individual, with the weights derived from the attention scores of each channel. At the regional and regional interaction level, we average the integrated attention matrices across all individuals to derive a matrix representing the importance of interactions between regions. At the regional level, we further compute the row-wise mean of this matrix to obtain the final regional importance scores. We visualize the top ten ROIs and interactions of ROIs with the highest attention scores using BrainNet Viewer (Xia et al., 2013), respectively.
@@
Alzheimer’s disease (AD). From the Fig. 12, we observed a significant overlap between the ROIs and those implicated in critical interactions, particularly involving the left caudate nucleus. This may be attributed to the damage sustained by relevant brain regions during the progression of Alzheimer’s disease, which subsequently leads to weakened connectivity between brain regions. The ten ROIs with the highest attention scores include: Right Caudate (CAU.R), Left Hippocampus (HIP.L), Right Parahippocampal Gyrus (PHG.R), Left Amygdala (AMYG.L), Left Olfactory Cortex (OLF.L), Left Middle Temporal Gyrus (TPOmid.L), Right Olfactory Cortex (OLF.R), Left Parahippocampal Gyrus (PHG.L), Left Caudate (CAU.L), and Right Heschl’s Gyrus (HES.R). These regions are primarily associated with functions such as memory (Hippocampus, Parahippocampal gyrus), emotion (Amygdala), olfaction (Olfactory cortex), motor control (Caudate nucleus), and auditory processing (Heschl’s gyrus). Specifically, the Hippocampus and Parahippocampal gyrus have been widely confirmed to be associated with AD (Bv and Agrawal, 2022; Echávarri et al., 2010), as they are involved in brain memory functions. The Amygdala has also been widely demonstrated to undergo atrophy during the progression of diseases (Zidan et al., 2019). And the Caudate, due to its crucial role in regulating motor control, has significant associations with AD (Zhi et al., 2024). Ye et al. also confirmed the association between the Middle Temporal Gyrus and AD (Ye et al., 2019). The Olfactory Cortex and the Heschl’s Gyrus currently lack direct evidence linking them to Alzheimer’s disease. However, they may be associated with the loss of hearing and smell in Alzheimer’s patients, which suggests that this finding obtained by HOI-brain may provide new insights into the pathology of Alzheimer’s disease. In addition, we observed that most of the top ten ROI interactions with the highest attention scores revolve around the Right Caudate, suggesting that the Right Caudate may serve as a hub node in the brain network and is disrupted during the progression of Alzheimer’s disease, which is consistent with the literature (Zhi et al., 2024).
@@
Parkinson’s syndrome (PD). As illustrated in Fig. 13, the ten ROIs with the highest attention scores are: Right Precentral Gyrus (PCG.R), Left Medial Orbitofrontal Cortex (ORBsupmed.L), Right Inferior Parietal Lobule (IPL.R), Left Amygdala (AMYG.L), Right Triangular Part of the Inferior Frontal Gyrus (IFGtriang.R), Right Superior Parietal Gyrus (SPG.R), Right Inferior Orbitofrontal Cortex (ORBinf.R), Right Thalamus (THA.R), Right Superior Temporal Gyrus (STG.R), and Right Opercular Part of the Inferior Frontal Gyrus (IFGoperc.R). These regions are predominantly implicated in motor execution, reward processing, executive control, affective regulation, and sensorimotor integration—domains known to be disrupted in PD. Specifically, the Precentral Gyrus, a key component of the primary motor cortex, exhibits marked hypometabolism and altered functional connectivity in PD (Wang et al., 2024b). The left medial orbitofrontal cortex is related to decision-making performance in Parkinson’s disease (Kobayakawa et al., 2016). Moreover, compared with normal controls, patients with PD showed a globally reduced structural-functional connectivity decoupling in the Inferior and Medial Orbitofrontal Cortices (ORBinf.R and ORBsupmed.L) (Zou et al., 2024). The Amygdala, central to emotional processing and reward evaluation, demonstrates reduced volume and disrupted connectivity within limbic-striatal circuits in PD, which may underlie affective dysregulation and apathy (Yoshimura et al., 2005). The Thalamus, a vital relay hub within cortico-basal ganglia-thalamo-cortical loops, is frequently implicated in PD pathophysiology due to its role in motor and non-motor symptomatology, including tremor and cognitive fluctuations (Zirh et al., 1998). The Triangular and Opercular parts of the Inferior Frontal Gyrus (IFGtriang.R and IFGoperc.R), critical for executive control and language processing, demonstrate impaired activation and connectivity in PD, aligning with cognitive inflexibility and verbal fluency deficits (Li et al., 2020). It is worth noting that although there is currently no direct literature indicating a pathological association between the superior temporal gyrus and PD, its involvement in auditory processing and social cognition may indirectly contribute to communication difficulties and social withdrawal symptoms in patients with Parkinson’s disease. Crucially, unlike ADNI and ASD, the ”core hubs” in the interaction network—PCG.R, STG.R, IFGtriang.R, and ANG.R—did not rank among the top positions in the ”key brain importance” list. Conversely, the seven regions listed as ”key brain importance,” such as ACG.R, TPOmid.R, and CAU.L, were almost entirely absent from the top 20 strongest interactions. Together, these findings highlight two distinct perspectives on ”importance” in brain: ”global centrality” versus ”local strong coupling.”

### v4

@@ 5.2 Model Comparison
Tables 2–5 present the comparative results across the ADNI, TaoWu, PPMI, and ABIDE datasets. Performance metrics (mean ± standard deviation) from ten-fold cross-validation are reported, with the best and second-best results highlighted in bold and underlined, respectively.
Several key observations emerge from our comparative analysis. First, general-purpose GNN-based models demonstrate comparable performance to traditional machine learning methods in most cases, without showing significant advantages. This can be attributed to neural networks’ increased parameter count when utilizing edge features, rendering them more prone to overfitting while maintaining better interpretability. Second, Transformer-based models consistently outperform GNN-based approaches across all four datasets, suggesting their superior capacity for capturing global connectome structures. Third, HGNN-based models show limited performance advantages, with notable results only on the PPMI dataset. We attribute this to: (1) their excessive parameterization increasing overfitting risk, and (2) their k-hop neighbor modeling failing to adequately represent co-fluctuation patterns among brain regions, thus poorly capturing HOIs. Notably, PH-based models (e.g., ATPGCN and Brain-HORS) demonstrate significant superiority over both GNN-based and traditional machine learning approaches by incorporating persistent homology-derived higher-order loop and connected component features. This underscores the critical importance of higher-order topological features in brain disease diagnosis.
@@
Our HOI-brain outperforms all 20 baselines on all datasets. In particular, it can be found that the proposed HOI-brain outperforms the traditional machine learning models on all datasets, with average accuracy improved by 12.5 (%), average precision improved by 12.9(%), average recall improved by 14.3(%) and F1-score improved by 12.7(%), respectively. And the proposed HOI-brain outperforms the GNN-based models on all datasets, with average accuracy improved by 19.6(%), average precision improved by 22.9(%), average recall improved by 19.5(%), and F1-score improved by 19.5(%), respectively. This indicates that our model, by incorporating higher-order topological features, is more beneficial for brain disease diagnosis compared to using only low-order edge features. Furthermore, the proposed HOI-Brain outperforms Transformer-based models on all datasets, with average accuracy improved by 13.0(%), average precision improved by 13.4(%), average recall improved by 10.3(%), and F1-score improved by 10.9(%), respectively. This highlights that our model, through the multi-channel brain network Transformer, can effectively and holistically integrate complementary information from heterogeneous higher-order and low-order topological features while preserving channel-specific properties. Next, when compared to HGNN-based models, HOI-Brain achieves remarkable improvements with average accuracy improved by 21.8(%), precision improved by 27.0(%), recall improved by 19.4(%), and F1-score improved by 20.5(%). These results clearly demonstrate that our proposed HOI-Brain framework significantly advances the modeling of HOIs through the Multiplication of Temporal Derivatives and the extraction of topological features using Persistent Homology in brain networks. When specifically compared to Persistent Homology (PH) based models, HOI-Brain achieves significant improvements with average accuracy increased by 17.4(%), precision enhanced by 21.4(%), recall boosted by 14.5(%), and F1-score elevated by 16.2(%). This is partly attributed to our method linking signed HOIs in the brain to higher-order organizations, and partly due to our exploration of more complex signed quadruplet structures and 2D voids formed by triangular faces and quadruplet.
@@ 5.3.1 Ablation Study
To evaluate the rationality and effectiveness of the proposed model’s architecture, we conducted a series of ablation experiments to assess the impact of different components on overall performance. Our first aim was to evaluate the effectiveness of more complex signed quadriplet structures and 2D voids formed by triangular faces. Specifically, following the method in (Santoro et al., 2024), we extracted lower-order edge features, violating triangles, and one-dimensional loops. In addition, we extracted unsigned good quadruplets and unsigned two-dimensional voids based on our methods. We implement it within different combinations of these features on all datasets. The experimental results presented in Tables 6, 7, 8, 9, indicate that integrating both higher-order and lower-order interactions proves to be more effective for brain disorder diagnosis compared to using lower-order interactions alone. Furthermore, quadruplet interactions captured by HOI-Brain have been demonstrated to be more effective in improving diagnosis than triplet interactions. This may be because a significant proportion of triplet interactions can be decomposed into a linear combination of pairwise interactions, which is insufficient to capture the true HOIs. Additionally, distinguishing between positively and negatively synergistic HOIs to generate signed higher-order structures, as opposed to unsigned ones, can further improve the effectiveness of brain disease diagnosis, which may be due to the fact that it can provide detailed insights into the communication within the brain.
Our second aim was to evaluate the effectiveness of different components in multi-channel brain network Transformer. The different variants are: not adding signed higher-order features decoupling mechanism (wo-signed), not adding attention-guided feature fusion mechanism (wo-fusion), and not adding orthonormal clustering readout of multiple channels(wo-cluster). From Figure 5, we can see that adding signed higher-order features decoupling mechanism and attention-guided feature fusion mechanism are always helpful as HOI-brain outperforms all variants, including (wo-signed) and (wo-fusion). This result is consistent with our assumption that by introducing the attention mechanism, our model can adaptively fuse both positive/negative information and multi-channel information, thereby emphasizing discriminative features while suppressing redundant information. Incorporating orthonormal clustering readout across multiple channels (wo-cluster) is highly beneficial, as this approach effectively captures both lower-order and higher-order feature patterns, which often reside within distinct functional modular structures.
@@ 5.3.2 Hyperparameter Analysis
To further demonstrate how the design of orthonormal clustering readout across multiple channels affects our model’s performance, we investigate a key hyperparameter: the number of clusters K. Specifically, we evaluate the method with K set to 2, 3, 4, 5, 10, 20, 50, and 100. The accuracy metric results when tuning this hyperparameter across all datasets are presented in Figure 6. The results of other metrics are presented in Figure 19. We observe that the model’s performance improves as K increases from 2 to 10 or 20 but declines when K rises from 10 or 20 to 100. This suggests that the optimal number of clusters is relatively small, reducing computational cost while aligning with the fact that the typical number of functional modules is fewer than 25, a finding consistent with previous studies (Kan et al., 2022b).
@@ 5.4.1 In-depth Analysis of Attention Mechanism
To better understand the mechanism of our model, which focuses on the most discriminative information in the brain for the diagnosis of brain diseases, we first performed a visualization analysis of the attention scores , of the attention-guided feature fusion mechanism for all patients of each type of disease, using ADNI and ABIDE as examples. Figure 7 shows that our model is capable of adaptively learning weights according to the characteristics of different diseases, thereby integrating higher-order and lower-order topological information for the diagnosis of brain diseases. In addition, we observed that quadruplet-level interaction signatures are significantly more important than two-dimensional void descriptors and edge features, with their average attention scores exceeding 0.5 on both datasets.
@@
Next, we also conducted a visualization analysis of the attention scores , , , derived from the signed higher-order feature decoupling mechanism for all patients across each disease type, using ADNI and ABIDE as examples. Figure 8 shows that across both datasets, the importance of negatively synergistic quadruplets is generally greater than that of positively synergistic information in most cases, while the importance of positively synergistic voids exceeds that of negatively synergistic information. This new phenomenon may provide a clearer direction for research into the pathological mechanisms of brain diseases.
@@
In order to conduct a more in-depth investigation into the aforementioned phenomena in the attention mechanism, inter-group comparisons were conducted of the signed higher-order topological features, the quadruplet-level interaction signatures and two-dimensional void descriptors. These were captured by the model in both the patient and healthy control groups. The ADNI dataset was used as an example. As demonstrated in Figure 9, which presents a comparison of group differences in the number of higher-order topological features, it was observed that the group differences in the quadruplet-level interaction signatures were found to be highly significant, with all statistical significance levels at compared to void descriptors. This finding is consistent with the results of our model visualisation. This phenomenon may be attributed to the disruption of HOIs within the brain during the progression of the disease, which in turn exerts an influence on the high-dimensional hole structures present within the brain. Furthermore, as the disease progresses, the number of positive quadruplets exhibits a gradual decrease, while the number of negative quadruplets initially decreases and subsequently increases. This finding suggests that during the transition from CN to MCI, there is a decline in synergistic interactions in the brain, whether positive or negative, as brain communication becomes progressively disrupted. Conversely, as the condition progresses from MCI to AD, brain regions have been observed to exhibit more negatively synergistic interactions, concurrently reducing their excitability to preserve the functionality of diverse neural modules. This finding aligns with the reference stating that excitatory neurons in the brains of Alzheimer’s patients suffer severe genomic damage (Miller et al., 2022). With regard to the number of voids, although no clear trend was observed in the progression of the disease, it was found that positive voids exhibited more significant inter-group differences compared to negative voids. This finding lends further support to the hypothesis that they exhibit higher levels of attention scores. Consequently, CN appears to utilise positive information to construct the whole-brain topological structure, whereas AD relies more heavily on negative information. These findings may offer a novel perspective on the pathological mechanisms underlying Alzheimer’s disease.
@@ 5.4.2 In-depth Analysis of Cluster Mechanism
To better understand how our model leverages modular-level similarities between ROIs across different brain network patterns, we first provide a visualization of three attention matrices at the first layer of the multi-head self-attention module on the ADNI and ABIDE datasets in Figures 10, 11. The attention scores are averaged across all subjects in the ADNI or ABIDE test set. This figure demonstrates that the learned attention scores reveal varying degrees of modularity in both lower-order and higher-order patterns, highlighting the interpretability of our model. Furthermore, we provide a visualization of three cluster assignment matrices on the ADNI or ABIDE dataset in Figure 12. The cluster assignment matrices are multiplied across layers and averaged across all subjects in the ADNI and ABIDE test set. Each row corresponds to an ROI and each column is a cluster. We observe that the assignment matrix, which is calculated based on the similarity between nodes and clusters, contains only one high-value entry per row in most cases, despite the fact that we did not impose any regularization constraints on the assignment matrix. This indicates that our model can effectively partition nodes into functional modules. Additionally, there are differences between the assignment matrices of low-order and higher-order patterns, which further demonstrates that varying degrees of modularity in both low-order and higher-order patterns. This result aligns with the visualization of three attention matrices, further indicating the interpretability of our model.
@@ 5.4.3 The Important Brain Regions and Interactions of Brain Regions
In this section, we further investigate whether the learned ROIs attention scores are interpretable and consistent with previous findings. Here, we take the ADNI, PPMI and ABIDE datasets as examples for discussion of three diseases. These three datasets contain larger sample sizes, enabling the reflection of more generalizable phenomena for brain diseases. First, we analyze the important brain regions and interactions of brain regions associated with brain diseases using HOI-Brain from a more comprehensive perspective rather than a low-level perspective as in previous studies (Kan et al., 2022b; Yu et al., 2024). For each subject, HOI-brain generates three attention matrices at the first layer of multi-head self attention module. Higher attention scores indicate greater discriminative power of the regions for AD, PD, and ASD. For each disease, we obtain a comprehensive attention matrix that simultaneously captures both lower-order and higher-order patterns by performing a weighted summation of the three attention matrices for each individual, with the weights derived from the attention scores of each channel. At the regional and regional interaction level, we average the integrated attention matrices across all individuals to construct a matrix representing the importance of interactions between regions. At the regional level, we further compute the row-wise mean of this matrix to obtain the final regional importance scores. We visualize the top ten ROIs and interactions of ROIs with the highest attention scores using BrainNet Viewer (Xia et al., 2013), respectively.
@@
Alzheimer’s disease (AD). From the Figure 13, we observed a significant overlap between the ROIs and those implicated in critical interactions, particularly involving the left caudate nucleus. This may be attributed to the damage sustained by relevant brain regions during the progression of Alzheimer’s disease, which subsequently leads to weakened connectivity between brain regions. The ten ROIs with the highest attention scores include: Right Caudate (CAU.R), Left Hippocampus (HIP.L), Right Parahippocampal Gyrus (PHG.R), Left Amygdala (AMYG.L), Left Olfactory Cortex (OLF.L), Left Middle Temporal Gyrus (TPOmid.L), Right Olfactory Cortex (OLF.R), Left Parahippocampal Gyrus (PHG.L), Left Caudate (CAU.L), and Right Heschl’s Gyrus (HES.R). These regions are primarily associated with functions such as memory (Hippocampus, Parahippocampal gyrus), emotion (Amygdala), olfaction (Olfactory cortex), motor control (Caudate nucleus), and auditory processing (Heschl’s gyrus). Specifically, the Hippocampus and Parahippocampal gyrus have been widely confirmed to be associated with AD (Bv and Agrawal, 2022; Echávarri et al., 2010), as they are involved in brain memory functions. The Amygdala has also been widely demonstrated to undergo atrophy during the progression of diseases (Zidan et al., 2019). The Caudate, due to its crucial role in regulating motor control, has significant associations with AD (Zhi et al., 2024). Ye et al. also confirmed the association between the Middle Temporal Gyrus and AD (Ye et al., 2019). The Olfactory Cortex and the Heschl’s Gyrus currently lack direct evidence linking them to Alzheimer’s disease. However, they may be associated with the decline of hearing and smell in Alzheimer’s patients, which suggests that this finding obtained by HOI-brain may provide new insights into the pathology of Alzheimer’s disease. In addition, we observed that most of the top ten ROI interactions with the highest attention scores revolve around the Right Caudate, suggesting that the Right Caudate may serve as a hub node in the brain network and is disrupted during the progression of Alzheimer’s disease, which is consistent with the literature (Zhi et al., 2024).
@@
Parkinson’s disease (PD). As illustrated in Figure 14, the ten ROIs with the highest attention scores are: Right Precentral Gyrus (PCG.R), Left Medial Orbitofrontal Cortex (ORBsupmed.L), Right Inferior Parietal Lobule (IPL.R), Left Amygdala (AMYG.L), Right Triangular Part of the Inferior Frontal Gyrus (IFGtriang.R), Right Superior Parietal Gyrus (SPG.R), Right Inferior Orbitofrontal Cortex (ORBinf.R), Right Thalamus (THA.R), Right Superior Temporal Gyrus (STG.R), and Right Opercular Part of the Inferior Frontal Gyrus (IFGoperc.R). These regions are predominantly implicated in motor execution, reward processing, executive control, affective regulation, and sensorimotor integration—domains known to be disrupted in PD. Specifically, the Precentral Gyrus, a key component of the primary motor cortex, exhibits marked hypometabolism and altered functional connectivity in PD (Wang et al., 2024b). The left medial orbitofrontal cortex is related to decision-making performance in Parkinson’s disease (Kobayakawa et al., 2016). Moreover, compared with normal controls, patients with PD showed a globally reduced structural-functional connectivity decoupling in the Inferior and Medial Orbitofrontal Cortices (ORBinf.R and ORBsupmed.L) (Zou et al., 2024). The Amygdala, central to emotional processing and reward evaluation, demonstrates reduced volume and disrupted connectivity within limbic-striatal circuits in PD, which may underlie affective dysregulation and apathy (Yoshimura et al., 2005). The Thalamus, a vital relay hub within cortico-basal ganglia-thalamo-cortical loops, is frequently implicated in PD pathophysiology due to its role in motor and non-motor symptomatology, including tremor and cognitive fluctuations (Zirh et al., 1998). The Triangular and Opercular parts of the Inferior Frontal Gyrus (IFGtriang.R and IFGoperc.R), critical for executive control and language processing, demonstrate impaired activation and connectivity in PD, aligning with cognitive inflexibility and verbal fluency deficits (Li et al., 2020). It is worth noting that although there is currently no direct literature indicating a pathological association between the superior temporal gyrus and PD, its involvement in auditory processing and social cognition may indirectly contribute to communication difficulties and social withdrawal symptoms in patients with Parkinson’s disease. Importantly, unlike ADNI and ASD, the core hubs in the interaction network—PCG.R, STG.R, IFGtriang.R, and ANG.R—were not among the most highly ranked regions in terms of brain importance. Conversely, the seven brain regions identified as key in terms of importance—such as ACG.R, TPOmid.R, and CAU.L—were almost entirely absent from the top 20 regions with the strongest interactions. Together, these findings underscore two distinct interpretations of importance in the brain: global centrality versus local strong coupling.